# Towards Generalist Robot Learning from Internet Video: A Survey

### v3

@@ 4 Methods
4.1 Video Foundation Models
Foundation models are large deep learning models trained on large, diverse datasets. They maintain general knowledge and capabilities useful in a wide range of downstream settings. This has most notably been demonstrated in natural language processing with large language models (LLMs) (?, ?). In this section, we focus on video foundation model techniques, which can act as a general means for extracting robotics-relevant knowledge (e.g., knowledge regarding physics and behaviours) from large-scale internet video data. We are interested in these models and techniques in LfV due to the following use-cases (see Figure 4):
  • Using pretrained video foundation models: A pretrained video foundation model can be adapted for downstream robot applications. For example, a video prediction model can be adapted into a robot dynamics model (?).
  • Using video foundation model techniques and datasets: Techniques and datasets used originally for video foundation models can be customized for robotics purposes. For example, a robot foundation model can be trained on both video and robot data, using techniques inspired by video foundation modelling (?).
@@ 4.1 Video Foundation Models
We loosely categorise the video foundation model literature into video encoding methods (Section 4.1.1), video prediction methods (Section 4.1.2), and video-to-text generation methods (Section 4.1.3). In each subsection we give a very brief overview of different approaches after initially explaining why each method is relevant to LfV.
@@ 4.1.1 Video Encoders
Foundational video encoders ($p(z|\tau)$, where $z$ is a representation and $\tau$ is a video clip) can provide rich and robust video representations for downstream robotic applications. Specifically, they can useful as follows.
@@ Methods.
  • SOTA Approaches. The literature has proposed a number of approaches for learning rich spatio-temporal representations from video in a scalable manner. Video-text losses leverage text annotations to learn semantic representations from video. This can include: video-text contrastive losses (?, ?, ?, ?, ?), video-text matching (?), masked language modelling (?), and video-to-text losses (?, ?). Auto-encoding approaches have also proven popular. State-of-the-art methods have used masked auto-encoding (MAE) (?, ?, ?) and vector-quantized auto-encoding (VQ-AE) (?, ?, ?, ?, ?). A number of works have explored the use of student-teacher distillation losses (?, ?, ?). Joint-embedding prediction architectures (JEPA) have been proposed to better mitigate issues related to video’s high-dimensionality and noise in video (?).
@@
  • Conditioning information can simplify the video prediction problem and allow for more control over generated videos. Such conditioning is valuable for downstream robotics: it can allow us to simulate the effects of different ‘action’ strategies. ? (?) outline various popular conditioning schemes for video generation models $p(\tau|c)$. In particular, language conditioning can allow for flexible and intuitive control over generated video, at varying levels of detail (?, ?, ?). Amongst other possibilities, video generation can also be conditioned on future images (i.e, video in-filling, see ?) or on action information (?). We outline various action representations that may be suitable for conditioning video predictions in Section 4.2.1.
@@ 4.1.3 Video-to-Text Models
This section details to models with video-to-text capabilities. A capable video-to-text model can perform, for example, video question-answering or video summarization. A capable video-to-text foundation model could be valuable to robotics in the following ways.
  • High-quality representations: A capable video-to-text model will have robust, high-quality video representations. Compared to a video-only model, it will likely have improved high-level semantic representations. Compared to an image-to-text model, it will have improved temporal-dynamic representations. Robotics can bootstrap from such models via representation transfer (?), or by adding robot data directly into the model’s pretraining corpus (?).
@@ Methods.
We now give an overview of existing video-to-text methods and models. We note that research here is somewhat preliminary: low-quality video captions and the difficulties of the video modality mean progress lags behind foundation models in other domains.
  • Compositional approaches have often been used to obtain video-to-text pipelines (?, ?, ?, ?). For example, ? (?) use an image-language model to answer questions about individual video frames, and an LLM to synthesize this information to produce a global summary of the video. However, the lack of end-to-end video training in compositional approaches mean they can lack rich, nuanced video representations.
  • Leveraging pretrained LLMs via adaptors and finetuning. Approaches here typically first obtain a pretrained LLM and an (often) pretrained video encoder, and then define an adaptor module to channel information from the video encoder output into the LLM (?, ?, ?, ?). Once the new video-to-text architecture is defined, a common training scheme is: (i) convert large, diverse video-text data into token sequences and perform next-token-prediction pretraining, then (ii) perform supervised instruction-tuning on small, high-quality instruction datasets (?, ?). Following trends in LLMs, future work may investigate a third RLHF finetuning stage (?). During training the LLM and video encoder may be finetuned (?), or kept frozen (?).
@@ 4.1.4 Challenges
Scaling of datasets and model sizes has recently lead to impressive advances in video ML (?, ?, ?). Most notably, the recent Sora model (?) demonstrated significant improvements in video prediction visual quality, physical realism, and generation lengths. However, there are still major limitations in the capabilities of current video foundation models. Models can hallucinate, and can struggle with spatial relationships, fine-grained spatio-temporal details, and long-term understandings and generation (?, ?, ?).
One bottleneck to further progress here is the quality (and quantity) of available video-data (see Section 5.1). In particular, improved annotations with precise low-level details will enhance robotics-relevant fine-grained understandings. The computational demands of processing high-dimensional and long video data present another issue. Improved model architectures can help here (?). For example, progress in long video understanding will be aided by more efficient architectures for handling longer contexts (?, ?, ?). Other directions to improve video foundation models include using RL finetuning to address hallucination issues (?), and leveraging 3D information to improve the physical realism of video generations (?). Improved methods for fine-grained, single-step conditioning of video predictions will be relevant to robotics (?). Finally, we advocate for progress in open-sourced foundational video prediction models: open-sourced models make LfV research more accessible to the wider community.
@@ 4.2.1 Action Representations
Here, we consider learned representations where $\hat{a}_{t}$ (i.e., the latent action) only contains information regarding the action that took us from $o_{t}$ to $o_{t+1}$. Such a latent representation is commonly learned using a next-observation prediction objective $p(o_{t+1}|\hat{a}_{t},o_{t})$, such that $\hat{a}_{t}$ is informative for the predictions of a forward dynamics model (FDM) (?, ?, ?, ?). A trainable (latent) inverse dynamics model (IDM) can be used to encode past-frames and infer $\hat{a}_{t}$ (?, ?). Here, some form of regularisation, such as a vector-quantization bottleneck (?, ?), can be employed to prevent the IDM from copying the entire observation $o_{t+1}$ into the latent action.
@@ 4.2.2 Representations to Address LfV Distribution-Shift
Distribution shift between internet video and the target robot domain poses a challenge to LfV approaches (see Section 3.4). For example, there exist embodiment differences between humans in internet video and robots. Such distribution shifts can hinder our ability to transfer knowledge from the video data to the robot.
A line of LfV research has attempted to overcome these issues by explicitly designing representations of (human) video that are more transferable to the robot domain. Once the transferable representation is obtained, works have most commonly used it to: (i) define a reward function encouraging the robot to match the behaviour in the video (?), or (ii) to help train a policy via behaviour cloning of the represented video (?).
@@
  • What types of human-embodiment information can be detected?
    (1) Poses: Several works explicitly estimate the pose of the human body or hand in videos (?, ?). This involves estimating the positions and orientations of various joints or key points on the body via off-the-shelf models. Once a human-hand pose is detected, it may need to be retargeted to the robot embodiment. This can be achieved by directly optimising a loss function at inference (?), or by training a retargeting network to minimise a retargeting loss function (?).
    (2) Affordances: A common affordance seen in the LfV literature is the combined use of grasp points and grasp trajectories (?, ?) – this abstraction transfers naturally from human to robot embodiment. This affordance information can be extracted from videos using an off-the-shelf model (?), and using further tricks to ensure accurate, usable affordance labels (?). Note, 2D affordances can be converted to 3D using depth estimation (?).
    (3) Masks and Bounding boxes: Masks or bounding boxes of human hands and objects can be used (?). These can be obtained from videos using off-the shelf models (?, ?) or via labelled datasets (?).
@@ Discussion.
Future directions: (1) An obvious route to improved performances here is to continue to scale our video dataset and model sizes (?, ?). Experimenting with scalable alternative-action techniques (?) will be useful for conditioning foundational video predictors. (2) A hierarchical world/dynamics model (?), where the higher levels are learned from video data and the lower-levels are learned from robot data, is another direction to explore. (3) An underexplored direction is to combine standard analytic simulation (?, ?) with video prediction simulation (?) to obtain the benefits of both. (4) Following synthetic data trends in other fields (?, ?), future work may seek to use foundational video predictors to generate synthetic video rollouts (?, ?), or augmenting existing data via video editing (?).
@@ 5.1.2 Curating Video Data
Diverse, large-scale video datasets can be obtained from internet repositories of pre-existing video data. Techniques here focus on ensuring only relevant and high-quality videos are scraped from these repositories. A typical pipeline for scraping internet video involves: (1) Formulating a pool of query prompts used to search the repository for candidate videos. Previous works have constructed this pool, for example, via: surveys of human time-use (?, ?); crowd-sourcing (?, ?); or by using LLMs to parse action prompts from text corpora (?) (2) Post-processing the pool of candidate videos. This has included processing videos into clips which contain no ‘cuts’ between scenes (?) and using video meta data (e.g., view counts) to identify and remove low-quality video (?, ?, ?). Automated metrics such as optical flow (?), image-language model embeddings (?, ?), and video ‘realism’ estimates (?) have also been used to filter video data. (3) Re-annotating the videos to improve the quality of labels. We describe video captioning methods in detail in the below paragraphs.
@@ 5.2 Benchmarks
5.2.1 Designing Benchmarks for LfV
In the paragraphs below, we first give recommendations regarding the metrics an LfV benchmark should measure, before outlining different possible categories of LfV benchmark and their corresponding desiderata.
@@ 6.1 High-level Recommendations
After our analysis of the LfV literature in Section 4, we now provide some high-level recommendations for future LfV research.
@@
We should focus on methods that can scale well to large, diverse internet video. Many previous LfV works limit scalability by making strong assumptions on the nature of the video data or the downstream robot setting (?, ?). We advocate for methods that can use general learning objectives to extract knowledge from heterogenous video data (see the foundation model approaches in Section 4.1). Methods that can learn from unlabelled and suboptimal video data are most scalable in theory (?), though methods that leverage language labels can still be highly effective (?, ?). Finally, we should focus initially on LfV approaches that can learn from offline robot data (or that can perform online RL in simulation), due to the impracticalities of online real-world RL.
@@
Future LfV research should focus more on obtaining the most promising LfV benefits (Section 3.3); namely, improving robot data-efficiency and improving generalization beyond $\mathcal{D}_{\text{robot}}$. Much past LfV research has sought to extract reward functions from video (Section 4.3.3). Whilst these approaches are useful, we believe learning policies (Section 4.3.1) and dynamics models (Section 4.3.2) is most promising for reducing reliance on robot data (as is discussed in their respective sections).
@@
Improving and scaling our LfV video datasets is a reliable way to improve our LfV models. In Section 5.1.3, we recommend scraping larger-scale video data from the internet, and captioning this data with high-quality language annotations. We advocate for the open-sourcing of such curated video datasets.
@@
We should explicitly measure the extent to which a method can handle LfV challenges and can provide the potential benefits of LfV. In Section 5.2.1, we outline concrete metrics for this. It is often difficult to identify these metrics in existing LfV research. In Section 5.2.2, we advocate for the design of improved LfV benchmarks that can offer these metrics off-the-shelf.
Combine LfV with complementary and alternative approaches.
Any practical LfV approach should seek to obtain as useful a $\mathcal{D}_{\text{robot}}$ as possible. In some cases, video data could aid the collection of (real and simulated) robot data. For example, video data and video models could help guide automatic generation of diverse and realistic simulated environments (?, ?). Elsewhere, we note that the utility of video data is not restricted to monolithic foundation model approaches or pure learning approaches. The components of a modular approach—components such as specialist agents (?) or vision systems—may benefit from the use of internet video data. Meanwhile, if learning is combined with more classical or structured control methods (?, ?, ?), video data could benefit the learning aspect.
@@ 6.2 Video Foundation Models for LfV
A highly promising LfV direction is to utilise large-scale internet video data to help train large foundation models (FMs) for robotics. Here, we outline some related research avenues and discussion points.
@@
Advancing video FMs and techniques will be a key driver of progress for scalable LfV approaches (Section 4.1). Key issues to tackle here include addressing dataset limitations and prohibitive computational requirements. More details on avenues for improving video FMs are discussed in Section 4.1.4.
@@
Generic video foundation models can be adapted into a RL knowledge modality to aid robot learning. Strong baselines here will include finetuning the video FM into an action-outputting policy using offline robot data (?, ?), or (if the video FM has video generation capabilities) finetuning it into a robot dynamics model. Data-efficient and computationally efficient adaptation mechanisms (?, ?) may be useful here.
@@
The most effective LfV approaches will likely use pipelines fully optimized for robotics purposes. We now touch on directions for customising video FMs and techniques for LfV.
  • Improve the capabilities of generic video FMs in areas relevant to robotics. For example, we could improve the physics realism of generic video prediction models (e.g., via RL finetuning; see ?), or the fine-grained understandings of generic video-to-text models (e.g., via improved fine-grained language captioning of the video data). The inclusion of domain-specific robot videos in generic video FM pretraining could improve its performance in the downstream robot setting.

### v4

@@ 4 Methods
4.1 Extracting Knowledge from Internet Video
Internet video lacks action labels and has various distribution shifts with the robot domain. Given these LfV challenges (see Section 3.4), a key LfV question is: How to extract robotics-relevant knowledge from large-scale, heterogenous internet videos? The video foundation model (FM) literature offers potential answers here.
Video FMs are large deep learning models trained on large internet-scraped video datasets. From this data, they extract general knowledge and capabilities useful in a wide range of downstream settings (?). This includes robotics-relevant knowledge related to semantics, physics, and behaviours. As such, video FMs and corresponding techniques can be highly applicable for the following LfV use-cases (see Figure 4):
  • Using pretrained models: A pretrained video FM can be transferred or adapted (e.g., finetuned) for downstream robot applications, allowing the robot to leverage its broad world knowledge. For example, a video prediction model can be adapted into a robot dynamics model (?).
  • Using techniques and datasets: Techniques and datasets originally used for video FMs can be customized for robotics purposes. For example, a robot FM can be trained on both video and robot data, using techniques inspired by video FMs (?). These techniques can allow LfV to better extract representations and knowledge from internet video, thus improving the generality of the robot.
The promise of these video FM use-cases is backed up by prior work demonstrating that text and image FMs can benefit robotics. Scalable architectures transferred from these other domains have allowed robot techniques to better fit large, diverse robot data (?, ?). Knowledge and representations from text and image FMs have been transferred to the robot domain to improve capabilities, generalization, and common-sense (?, ?, ?). The use of video FMs may lead to further improvements due to the unique robotics-relevant information video data can offer.
Video foundation models and techniques will likely be key driver to the success of scalable LfV approaches. As such, this section provides LfV researchers with essential information regarding potential use-cases and the lower-level workings of these techniques. We loosely categorise the video foundation model literature into: video encoding methods (Section 4.1.1); video prediction methods (Section 4.1.2); and video-to-text generation methods (Section 4.1.3).
@@ 4.1.1 Video Encoders
Foundational video encoders ($p(z|\tau)$, where $z$ is a representation and $\tau$ is a video clip) can provide rich and robust video representations for downstream robotic applications. Specifically, they can useful as follows:
@@ Methods.
  • SOTA Approaches. The literature has proposed a number of approaches for learning rich spatio-temporal representations from video in a scalable manner. Video-text losses leverage text annotations to learn semantic representations from video. This can include: video-text contrastive losses (?, ?, ?, ?, ?), video-text matching (?), masked language modelling (?), and video-to-text losses (?, ?). Auto-encoding approaches have also proven popular. State-of-the-art methods have used masked auto-encoding (MAE) (?, ?, ?) and vector-quantized auto-encoding (VQ-AE) (?, ?, ?, ?, ?). A number of works have explored the use of student-teacher distillation losses (?, ?, ?). Joint-embedding prediction architectures (JEPA) have been proposed to mitigate issues related to video’s high-dimensionality and noise in video (?).
@@
  • Conditioning information can simplify the video prediction problem and allow for more control over generated videos. Such conditioning is valuable for downstream robotics: it can allow us to simulate the effects of different ‘action’ strategies. ? (?) outline various popular conditioning schemes for video generation models $p(\tau|c)$. In particular, language conditioning can allow for flexible and intuitive control over generated video, at varying levels of detail (?, ?, ?). Amongst other possibilities, video generation can also be conditioned on future images (i.e, video in-filling, see ?) or on action representations (?). We outline various action representations that may be suitable here in Section 4.2.1.
@@ 4.1.3 Video-to-Text Models
This section details to models with video-to-text capabilities, $p(\text{text}|\tau)$. A capable video-to-text model can perform, for example, video question-answering or video summarization. Such a video-to-text foundation model could be valuable to robotics in the following ways.
  • High-quality representations: A capable video-to-text model will have robust, high-quality video representations. Compared to a video-only model, it will have improved high-level semantic representations. Compared to an image-to-text model, it will have improved temporal-dynamic representations. Robotics can bootstrap from such models via representation transfer (?), or by adding robot data directly into the model’s pretraining corpus (?).
@@ Methods.
We now give an overview of existing video-to-text methods and models. We note that research here is somewhat preliminary: low-quality video captions and the difficulties of the video modality means progress lags behind foundation models in other domains.
  • Compositional approaches have often been used to obtain video-to-text pipelines (?, ?, ?, ?). For example, ? (?) use an image-language model to answer questions about individual video frames, and an LLM to synthesize this information to produce a global summary. However, the lack of end-to-end video training in compositional approaches mean they can lack rich, nuanced video representations.
  • Leveraging pretrained LLMs via adaptors and finetuning. Approaches here typically first obtain a pretrained LLM and an (often) pretrained video encoder, and then define an adaptor module to channel information from the video encoder output into the LLM (?, ?, ?, ?). Once the new video-to-text architecture is defined, a common training scheme is: (i) convert large, diverse video-text data into token sequences and perform next-token-prediction pretraining, then (ii) perform supervised instruction-tuning on small, high-quality instruction datasets (?, ?). Following trends in LLMs, future work could investigate a third RLHF finetuning stage (?). During training, the LLM and video encoder may be finetuned (?), or kept frozen (?).
@@ 4.1.4 Challenges
Scaling datasets and model sizes has recently lead to impressive advances in video ML (?, ?, ?). Most notably, the recent Sora model (?) demonstrated significant improvements in video prediction visual quality, physical realism, and generation lengths. However, there are still major limitations in the capabilities of current video foundation models. Models can hallucinate, and can struggle with spatial relationships, fine-grained spatio-temporal details, and long-term understandings and generation (?, ?, ?).
One bottleneck to further progress here is the quality (and quantity) of available video-data (see Section 5.1). In particular, improved annotations with precise low-level details will enhance robotics-relevant fine-grained understandings. The computational demands of processing high-dimensional and long video data present is another challenge. Improved model architectures can help here (?). For example, progress in long video understanding will be aided by efficient architectures for handling longer contexts (?, ?, ?). Other directions to improve video foundation models include using RL finetuning to address hallucination issues (?), and leveraging 3D information to improve the physical realism of video generations (?). Improved methods for fine-grained, single-step conditioning of video predictions will be relevant to robotics (?). Finally, we advocate for progress in open-sourced foundational video prediction models: open-sourced models make LfV research more accessible to the wider community.
@@ 4.2.1 Action Representations
Here, we consider learned representations where $\hat{a}_{t}$ (i.e., the latent action) only contains information regarding the action that took us from $o_{t}$ to $o_{t+1}$. Such a latent representation is commonly learned using a next-observation prediction objective $p(o_{t+1}|\hat{a}_{t},o_{t})$, such that $\hat{a}_{t}$ is informative for the predictions of a forward dynamics model (FDM) (?, ?, ?, ?, ?). A trainable (latent) inverse dynamics model (IDM) can be used to encode past-frames and infer $\hat{a}_{t}$ (?, ?). Here, some form of regularisation, such as a vector-quantization bottleneck (?, ?), can be employed to prevent the IDM from copying the entire observation $o_{t+1}$ into the latent action.
@@ 4.2.2 Representations to Address LfV Distribution-Shift
Distribution shift between internet video and the target robot domain poses a challenge to LfV approaches (see Section 3.4). Such distribution shifts can hinder our ability to transfer knowledge from the video data to the robot. A line of LfV research has attempted to overcome these issues by explicitly designing representations of (human) video that are more transferable to the robot domain. Once the transferable representation is obtained, works have most commonly used it to: (i) define a reward function encouraging the robot to match the behaviour in the video (?), or (ii) to help train a policy via behaviour cloning of the represented video (?).
@@
  • What types of human-embodiment information can be detected? (1) Poses: Several works explicitly estimate the pose of the human body or hand in videos (?, ?, ?). This involves estimating the positions and orientations of various joints or key points on the body via off-the-shelf models. Once a human-hand pose is detected, it may need to be retargeted to the robot embodiment. This can be achieved by directly optimising a loss function at inference (?), or by training a retargeting network to minimise a retargeting loss function (?). (2) Affordances: A common affordance seen in the LfV literature is the combined use of grasp points and grasp trajectories (?, ?) – this abstraction transfers naturally from human to robot embodiment. This affordance information can be extracted from videos using an off-the-shelf model (?), and using further tricks to ensure accurate, usable affordance labels (?). Note, 2D affordances can be converted to 3D using depth estimation (?). (3) Masks and Bounding boxes: Masks or bounding boxes of human hands and objects can be used (?). These can be obtained from videos using off-the shelf models (?, ?, ?) or via labelled datasets (?).
@@ Discussion.
Future directions: (1) An obvious route to improved performances here is to continue to scale our video dataset and model sizes (?, ?). Experimenting with scalable alternative-action techniques (?) will be useful for conditioning foundational video predictors. (2) A hierarchical world/dynamics model (?), where the higher levels are learned from video data and the lower-levels are learned from robot data, is another direction to explore. (3) An underexplored direction is to combine standard analytic simulation (?, ?) with video prediction simulation (?) to obtain the benefits of both. (4) Following synthetic data trends in other fields (?, ?), future work may seek to use foundational video predictors to generate synthetic video rollouts (?, ?), or augment and improve the coverage of existing data via video editing (?).
@@ 5.1.2 Curating Video Data
Diverse, large-scale video datasets can be obtained from internet repositories of pre-existing video data. Techniques here focus on ensuring only relevant and high-quality videos are scraped from these repositories. A typical pipeline for scraping internet video involves: (1) Formulating a pool of query prompts used to search the repository for candidate video with relevant content. Previous works have constructed this pool, for example, via: surveys of human time-use (?, ?); crowd-sourcing (?, ?); or by using LLMs to parse action prompts from text corpora (?) (2) Post-processing the pool of candidate videos. This has included processing videos into clips which contain no ‘cuts’ between scenes (?) and using video meta data (e.g., view counts) to identify and remove low-quality video (?, ?, ?). Automated metrics such as optical flow (?), image-language model embeddings (?, ?), and video ‘realism’ estimates (?) have also been used to filter video data. (3) Re-annotating the videos to improve the quality of labels. We describe video captioning methods in detail in the below paragraphs.
@@ 5.2 Benchmarks
5.2.1 Designing LfV Benchmarks
In the paragraphs below, we first give recommendations regarding the metrics an LfV benchmark should measure, before outlining different possible categories of LfV benchmarks and their corresponding desiderata.
@@ 6.1 High-level Recommendations
Following our the analysis of the LfV literature in Section 4, we now provide some high-level recommendations for future LfV research.
@@
Methods that can scale well to large, diverse internet video may be particularly effective. Many previous LfV works limit scalability by making strong assumptions on the nature of the video data or the downstream robot setting (?, ?). Instead, we advocate for methods that can use general learning objectives to extract knowledge from heterogenous video data (see the foundation model approaches in Section 4.1). Methods that can learn from unlabelled and suboptimal video data are most scalable in theory (?), though methods that leverage language labels can still be highly effective (?, ?). Finally, focusing initially on LfV approaches that can learn from offline robot data (or that can perform online RL in simulation) may be sensible, due to the impracticalities of online real-world RL.
@@
Future LfV research should focus on obtaining the most promising LfV benefits (Section 3.3); namely, improving robot data efficiency and improving generalization beyond $\mathcal{D}_{\text{robot}}$. Much past LfV research has sought to extract reward functions from video (Section 4.3.3). Whilst these approaches are useful, learning policies (Section 4.3.1) and dynamics models (Section 4.3.2) is likely most promising for reducing reliance on robot data (as is discussed in their respective sections).
@@
Improving and scaling video datasets is a reliable way to improve our LfV models. In Section 5.1.3, we recommend scraping larger-scale video data from the internet, and captioning this data with high-quality language annotations. We advocate for the open-sourcing of such curated video datasets.
Combine LfV with alternative approaches, like simulation.
Any practical LfV approach should seek to collect as useful a $\mathcal{D}_{\text{robot}}$ as possible. In some cases, video data could aid the collection of (real and simulated) robot data. Automatic generation of diverse and realistic simulated environments (?, ?) could be aided by video data (?) and video models (?). Here, video foundation models could be used to provide dense reward signals in diverse simulated environments (Section 4.3.3, ?). Elsewhere, we note that the use of video data is not restricted to monolithic foundation model approaches or pure end-to-end learning approaches. Components in modular approaches, such as specialist agents (?) or vision systems, may benefit from the use of internet video data. In hybrid approaches, where deep learning is combined with more classical or structured control methods (?, ?, ?), video data can benefit the deep learning component.
@@
Research should explicitly measure the extent to which a method can handle LfV challenges and can provide the potential benefits of LfV. In Section 5.2.1, we outline concrete metrics for this. It is often difficult to identify these metrics in existing LfV research. In Section 5.2.2, we advocate for the design of improved LfV benchmarks that can offer these metrics off-the-shelf.
@@ 6.2 Video Foundation Models for LfV
A promising LfV direction is to utilise large-scale internet video data to help train large foundation models (FMs) for robotics. Here, we outline related research avenues and key points of discussion.
@@
Advancing video FMs and techniques may be a key driver of progress for scalable LfV approaches (Section 4.1). Key issues to tackle here include addressing dataset limitations and prohibitive computational requirements. More details on avenues for improving video FMs are discussed in Section 4.1.4.
@@
Generic video foundation models can be adapted into RL knowledge modalities to aid robot learning. Strong baselines here will include finetuning the video FM into an action-outputting policy using offline robot data (?, ?), or (if the video FM has video generation capabilities) finetuning it into a robot dynamics model. Data-efficient and computationally efficient adaptation mechanisms (?, ?) may be useful here.
@@
The most effective LfV approaches will likely use pipelines fully optimized for robotics. We now touch on directions for customising video FMs and techniques for LfV.
  • Improve the capabilities of generic video FMs in areas relevant to robotics. We could, for example, improve the physics realism of generic video prediction models (e.g., via RL finetuning; see ?), or the fine-grained understandings of generic video-to-text models (e.g., via improved fine-grained language captioning of the video data). The inclusion of domain-specific robot videos in generic video FM pretraining could improve its performance in the downstream robot setting.